elimination/addition
a - 3b = 4
a + 5b = -4
with steps too

Answer by solver15(27)

multipy -1 to the equation a-3b=4
your equation will be
-a+3b=-4
then elimate -a with a
-a+3b=-4
a+5b=-4
-a and a is now elimated, so the equation will be
8b=-8, the divide both sides by 8
8b/8=-8/8
then b=-1
plug in -1 to the previous equation
a-3(-1)=4
a+3=4, the subtract 3 form both sides of the equation
-3 -3
the a=1, so your solution is (1,-1)